OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = −x − 2, find (f + g)(−1). −4 4 −6 −8

OpenStudy (anonymous):

f(-1) =? g(-1) = ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

4(-1) +1= -3 -(-1)-2=-1 -3-1=-4

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

oh its -4

OpenStudy (anonymous):

yup